# GaMeS: Mesh-Based Adapting and Modification of Gaussian Splatting

### v3

@@ Abstract
Recently, a range of neural network-based methods for image rendering have been introduced. One such widely-researched neural radiance field (NeRF) relies on a neural network to represent 3D scenes, allowing for realistic view synthesis from a small number of 2D images. However, most NeRF models are constrained by long training and inference times. In comparison, Gaussian Splatting (GS) is a novel, state-of-the-art technique for rendering points in a 3D scene by approximating their contribution to image pixels through Gaussian distributions, warranting fast training and swift, real-time rendering. A drawback of GS is the absence of a well-defined approach for its conditioning due to the necessity to condition several hundred thousand Gaussian components. To solve this, we introduce the Gaussian Mesh Splatting (GaMeS) model, which allows modification of Gaussian components in a similar way as meshes. We parameterize each Gaussian component by the vertices of the mesh face. Furthermore, our model needs mesh initialization on input or estimated mesh during training. We also define Gaussian splats solely based on their location on the mesh, allowing for automatic adjustments in position, scale, and rotation during animation. As a result, we obtain a real-time rendering of editable GS.
Gaussian Splatting, Mesh, NeRF, Rendering
Before modification   After modification
Figure 1: GaMeS produce a hybrid of Gaussian Splating and mesh representations. Therefore, GaMeS allows real-time modification and adaptation of Gaussian Splatting. The supplementary material provides more examples, including a video illustrating our results.
@@ 1 Introduction
Recently, we have observed the emergence of several promising methods for rendering unseen views of 3D objects and scenes using neural networks. For instance, Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) have rapidly grown in popularity within the computer vision and graphics communities (Gao et al., 2022) as they enable the creation of high-quality renders. Despite this interest and growing body of related research, the long training and inference time remains an unsolved challenge for NeRFs.
In contrast, latterly introduced Gaussian Splatting (GS) (Kerbl et al., 2023) offers swift training and real-time rendering capabilities. What is unique to this method is that it represents 3D objects using Gaussian distributions (i.e. Gaussians). Hence, it does not rely on any neural network. Consequently, Gaussians can be employed in a manner akin to manipulating 3D point clouds or meshes, allowing for actions like resizing or repositioning in 3D space. Nonetheless, practical challenges may arise when altering Gaussian positions, particularly in accurately tracking changes in the shape of Gaussian components, such as the ellipses. Moreover, scaling Gaussian components proves challenging when the object undergoes resizing, which is not an issue for classical meshes, as their triangle faces can be readily updated when vertex positions are adjusted.
The above constraints may be resolved by constructing Gaussian directly on the mesh, as shown by SuGaR (Guédon & Lepetit, 2023). Here, the authors introduced a regularization term in the Gaussian splat cost function to promote optimal alignment of the Gaussians with the scene’s surface. The SuGaR model uses signed distance functions (SDF) from the vanilla GS and minimizes the difference between this SDF and its actual value computed for the Gaussians. Moreover, the authors propose a method for extracting mesh from the GS model and introduce an optional refinement strategy that binds Gaussians to the surface of the mesh and jointly optimizes these Gaussians and the mesh through GS rendering.
@@
In our paper, we propose an alternative approach to such a task. Unlike SuGaR, we do not have two expensive stages dedicated to producing high-quality mesh. Instead, our Gaussian Mesh Splatting (GaMeS) uses existing mesh or approximates it during one-stage training. Our estimated mesh uses a set of triangle faces, which are not connected; therefore, we called it pseudo-mesh. Our pseudo-mesh is dedicated to modifying the GS model, not approximating objects’ surfaces. Implementing GaMeS involves positioning Gaussian components on the faces of the pseudo-mesh, ensuring proper alignment of these components with the mesh structure. Using our method, we can obtain comparable to state-of-the-art high-quality outcomes that can be attained for static scenes akin to the GS method. Moreover, any alterations to the mesh will automatically propagate updates to the corresponding Gaussian components, enabling real-time animation (see Fig. 1). Such an approach can also be applied in scenarios with preexisting mesh (see Fig. 2). Most importantly, we can train GS and then generate pseudo-mesh from Gaussian components to force a model to be editable. In such a strategy, our pseudo-mesh is not dedicated to describing the object’s surface but to editing the Gaussian Splatting model.
In summary, this work makes the following contributions:
@@
  • GaMeS work on the simple pseudo-mesh that can be estimated directly from Gaussian components obtained in a single training without an expensive preprocessing strategy.
  • Our method relies only upon essential vector operation. Consequently, we can render dynamic scenes at a similar pace to their static counterpart.
@@ 2 Related Works
Point-based Gaussian representations have found a large number of potential application scenarios, including substituting point-cloud data (Eckart et al., 2016), molecular structures modeling (Blinn, 1982), or shape reconstruction (Keselman & Hebert, 2023). These representations can also be utilized in applications involving shadow (Nulkar & Mueller, 2001) and cloud rendering (Man, 2006). In addition, a new technique of 3D Gaussian Splatting (3D-GS) was recently introduced (Kerbl et al., 2023). This method couples splatting methods with point-based rendering for real-time rendering speed. The rendering quality of 3D-GS is comparable to that of Mip-NeRF (Barron et al., 2021), one of the best multilayer perceptron-based renderers.
The GS surpasses NeRF in terms of both training and inference speed, distinguishing itself by not depending on a neural network for operation. Rather, GS stores essential information within its 3D Gaussian components, a feature that renders it well-suited for dynamic scene modeling (Wu et al., 2023). Additionally, integrating GS with a dedicated 3D computer graphics engine is a straightforward process (Kerbl et al., 2023). However, conditioning GS is a challenging task due to the large number of Gaussian components typically involved, which can reach hundreds of thousands.
GaussianAvatars (Qian et al., 2023), utilizes the local coordinate system to generate Gaussians corresponding to the mesh’s faces. This approach is specifically designed for avatars and assumes the availability of a realistic (external) model for mesh fitting. However, training both the mesh and GS simultaneously is not feasible. While these solutions offer certain advantages, they do not directly combine Gaussian components with the mesh. As a result, automatic adaptation of the Gaussian parameters to changing mesh is impossible.
@@
The more advanced method is proposed by SuGaR (Guédon & Lepetit, 2023), which introduces a regularization term in the GS cost function to encourage alignment between the Gaussians and the scene surface. SuGaR achieves this by using signed distance functions (SDF) and minimizing the difference between the SDF and the computed Gaussian values. Furthermore, the authors propose a method for extracting mesh from the GS model and introduce an optional refinement strategy that binds Gaussians to the mesh surface and jointly optimizes these Gaussians and the mesh through GS rendering.
Our work shows we can produce an editable GS model without expensive preprocessing stages like those used in SuGaR.
@@ 3 GaMeS : Mesh-Based Gaussian Splatting
Here, we delve into the details of the GaMeS model, commencing with the fundamental aspects of vanilla GS. Subsequently, we elucidate how we parameterize Gaussian distributions on the mesh faces. Then, we investigate extracting a mesh face from a single Gaussian component. Such face is used to reparametrize Gaussian components to be editable. Finally, we introduce the novel GaMeS approach.
Gaussian Splatting
The GS technique captures a 3D scene through an ensemble of 3D Gaussians, each defined by its position (mean), covariance matrix, opacity, and color. Additionally, spherical harmonics (SH) are employed to depict the color attributes. (Fridovich-Keil et al., 2022; Müller et al., 2022). The GS effectiveness is primarily attributed to the rendering process, which utilizes projections of Gaussian components. GS relies on a dense set of 3D Gaussian components with color and opacity:
@@
In practice, GS uses factorization $\Sigma=R^{T}SSR$. We can train GS with a flat Gaussian distribution by forcing $s_{1}=0$, where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$.
The parameters of Gaussian distributions undergo direct training through gradient optimization. To enhance adaptability to complex scenes, GS employs additional training strategies. Notably, significant Gaussian components are subdivided. If the update parameters are substantial, Gaussians components are duplicated, which also can be removed due to their low transparency. In addition, the GS training procedure is implemented in the CUDA kernel, which supports fast training and real-time rendering.
Figure 4: The left image presents the Gaussian component constructed on the face $V=\{{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}\}$. We overlap the centers of Gaussian and face and parameterize the covariance matrix by vertices $V$. The right image presents how our model uses previously estimated Gaussian to construct $k$ (in Fig. $k=3$) components in GS. GaMeS components are built by scaling and a translation on a mesh face.
@@
In GaMeS, we placed all Gaussian components onto the mesh surface. Let us examine a single triangle face with vertices:
$V=\{{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}\}\subset\mathbb{R}^{3}.$
We aim to parameterize the Gaussian components using vertices from the face $V$. We express the mean vector as a convex combination of vertices $V$, thus determining the Gaussian splats positions:
@@
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are trainable parameters such that $\alpha_{1}+\alpha_{2}+\alpha_{3}=1$. Through this parameterization, we consistently maintain the Gaussians positioned in the middle of the face $V$.
Figure 5: GaMeS also works when we do not have a mesh. In such a situation, we estimate pseudo-mesh obtaining a set of disconnected faces. Our pseudo-mesh is dedicated to modifying the GS model, not approximating objects’ surfaces.
Original Gaussian Create face Parameterized Gaussian
Figure 6: Each Gaussian corresponds to one face from the pseudo-mesh. The model parameterizes the Gaussian depending on the vertices, allowing it to be edited.
@@
The covariance matrix can be defined as an empirical covariance calculated from three points. However, such a solution is complex to combine with the optimization proposed in GS. Instead, the covariance is parameterized by factorization:
$\Sigma=R^{T}SSR,$
where $R$ is the rotation matrix and $S$ the scaling parameters. Here, we define rotation and scaling matrices to stay in the original framework. Let start from orthonormal vectors: ${\bf r}_{1},{\bf r}_{2},{\bf r}_{3}\in\mathbb{R}^{3}$, consisted into rotation matrix $R_{V}=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}].$ Where the first vector is defined by the normal vector:
@@
The last operation is obtained through orthonormalizing the vector with respect to the existing two vectors (a single step in the Gram–Schmidt process (Björck, 1994)):
@@
In consequence, we obtain a rotation matrix $R_{V}=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, which aligns with the triangular face. As scaling parameter $S$ we use:
$S_{V}=\mathrm{diag}(s_{1},s_{2},s_{3}),$
where $s_{1}=\varepsilon,$ $s_{2}=\|\mathrm{mean}({\bf v}_{1},{\bf v}_{2},{\bf v}_{3})-{\bf v}_{1}\|,$ and $s_{3}=\langle{\bf v}_{2},{\bf r}_{3}\rangle.$ The first scaling parameters correspond with the normal vector. Since, in our case, the Gaussians are positioned on the faces, aligning with the surface $s_{1}$ should be equal to zero. To avoid numerical problems, we fix this value as a small value (constant).
Original mesh Subdivide larger faces
Refer to caption Refer to caption
Figure 7: GaMeS use a fixed number of faces. In the case of a mesh containing faces of different sizes, finding such parameters causes problems. In such a case, we propose to divide large faces into smaller parts. The left side includes only the original mesh, while on the right, the sizable mesh faces of the brown pool surrounding have been subdivided into smaller ones.
Refer to caption Refer to caption
Figure 8: GaMeS parameterizes all Gaussian components by mesh vertices to lie directly on the mesh. Therefore, when we modify the mesh positions, we automatically adapt the components’ means and covariances. Consequently, we can see differences in the rendered image quality for the transforming when (a) only the Gaussian position is editable and (b) when the Gaussian rotation and scaling are editable.
Consequently, the covariance of Gaussian distribution positioned to face is given by:
$\Sigma_{V}=R_{V}^{T}S_{V}S_{V}R_{V},$
and correspond with the shape of a triangle $V$. Fig.  4 shows the process of determining Gaussians. Here the $\alpha_{j}^{i}$ refer to the $i$-th Gaussian $\mathcal{G}_{V}^{i}$ and the $j$-th vertex, respectively.
@@
This distribution method enables us to render Gaussian splats solely dependent on the position of the mesh triangles. Consequently, we can readily perform geometric transformations when dealing with Gaussians within a single (triangle) face. Triangle transformation will apply Gaussian transformation as well.
Distribution on pseudo-mesh faces
When we do not have mesh or any initialization, we can automatically produce pseudo-mesh. First, we train GS with flat Gaussian’s using $s_{1}=\epsilon$ in training. Then, we estimate one triangle face for each Gaussian component and parameterize Gaussian components by the vertices of the mesh face, see Fig. 5. After reparametrization, we have final GaMeS.
We use one pseudo-mesh face for each Gaussian component. Let us assume that we have a Gaussian component parameterized by mean ${\bf m}$, rotation matrix $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$ and scaling $S=\mathrm{dig}(\varepsilon,s_{2},s_{3}).$ We define three vertex of our pseudo-mesh face, see, Fig. 6
$V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}],$
where ${\bf v}_{1}={\bf m}$, ${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$.
Now we froze the vertex of the face $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$ and reparameterize the Gaussian component by defining $\hat{\bf m}$, $\hat{R}=[\hat{\bf r}_{1},\hat{\bf r}_{2},\hat{\bf r}_{3}]$ and $\hat{S}=\mathrm{dig}(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}).$ First, we put $\hat{\bf m}={\bf v}_{1}$. The first vertex of $\hat{R}$ is given by a normal vector:
@@
where $\times$ is the cross product. The second one is defined by $\hat{\bf r}_{2}=\frac{({\bf v}_{2}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\|}.$ The third one is obtained as a single step in the Gram–Schmidt process (Björck, 1994):
@@
Scaling parameters can also be easily calculated as $s_{1}=\varepsilon$, $\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|$ and $\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle$.
Consequently, the covariance of Gaussian distribution positioned to face is given by:
$\hat{\Sigma}_{V}=\hat{R}_{V}^{T}\hat{S}_{V}\hat{S}_{V}\hat{R}_{V},$
and correspond with the shape of a triangle $V$. For one face $V$, we define the corresponding Gaussian component:
@@
We obtain the Gaussian component parameterized by face mesh. We can use such a strategy GS with a flat Gaussian component.
GaMeS: Gaussian Splatting on Mesh
The essential step in our model involves positioning Gaussians on the 3D mesh of the scene or object in question. Here, we have to consider two crucial experimental scenarios, namely, with and without using mesh during the training. First, we assume that a mesh is being provided and exclusively train the Gaussian components’ centers, colors, and opacity. Here, numerous methods facilitate the generation of an initial mesh for objects, but our primary focus is not on this task. Instead, we emphasize addressing distinct issues that extend beyond the scope of initializing object meshes like EG3D (Chan et al., 2022), NGLOD (Takikawa et al., 2021b) or NeRFMeshing (Rakotosaona et al., 2023).
Second, we assume that we do not have a mesh. In the latter case, initialization of the mesh is required and achieved by utilizing our strategy with extracting pseudo-mesh from pretrained GS.
Additionally, we also show the initialization of the mesh using Faces Learned with an Articulated Model and Expressions (FLAME) (Li et al., 2017). The mesh initialization of the human face is not perfected fitted. We adjust the mesh vertices and Gaussian parameters throughout the training process.
Let us consider a mesh denoted as $\mathcal{M}=(\mathcal{V},\mathcal{E}),$ where $\mathcal{V}\subset\mathbb{R}^{3}$ represents vertices, and $\mathcal{E}\subset\mathbb{R}^{3}$ signifies edges. The mesh’s faces, denoted by =$\mathcal{F}_{\mathcal{M}}=\left\{\left({\bf v}^{V_{i}}_{1},{\bf v}^{V_{i}}_{2},{\bf v}^{V_{i}}_{3}\right)\right\}_{i}^{n}$ are described by sets of vertices. As previously detailed, such mesh serves as a parameterization for shaping Gaussian components.
For our experiments, we choose the number $k\in\mathbb{N}$ of Gaussian components per mesh face and define how densely they should be placed on its surface. For mesh containing $n\in\mathbb{N}$ faces, the final number of Gaussians is fixed and equal to $k\cdot n$.
When we do not have mesh, we train GS, and for each Gaussian, we estimate face. Then we obtain set $\mathcal{G}_{\mathcal{M}}$ of 3D parameterized Gaussians $\hat{\mathcal{G}}_{V}$ on pseudo-mesh $\hat{\mathcal{M}}$:
Mesh initialization
During training, we can modify mesh vertices and Gaussian parameters. We can also use external tools for mesh fitting (Takikawa et al., 2021a) and train only Gaussian components. In the experimental section, we show both of such scenarios.
In our model, we use a fixed number of Gaussian components per face. In the case of models with various face sizes, finding such parameters is not trivial. To solve such a problem, we propose to divide large faces into smaller parts, see Fig. 7.
PSNR $\uparrow$

### v4

@@ Abstract
Gaussian Splatting (GS) is a novel, state-of-the-art technique for rendering points in a 3D scene by approximating their contribution to image pixels through Gaussian distributions, warranting fast training and real-time rendering. The main drawback of GS is the absence of a well-defined approach for its conditioning due to the necessity of conditioning several hundred thousand Gaussian components. To solve this, we introduce the Gaussian Mesh Splatting (GaMeS) model, which allows modification of Gaussian components in a similar way as meshes. We parameterize each Gaussian component by the vertices of the mesh face. Furthermore, our model needs mesh initialization on input or estimated mesh during training. We also define Gaussian splats solely based on their location on the mesh, allowing for automatic adjustments in position, scale, and rotation during animation. As a result, we obtain a real-time rendering of editable GS.
@@ 1 Introduction
Figure 1: GaMeS produce a hybrid of Gaussian Splatting (GS) and mesh representations. Therefore, GaMeS allows real-time modification and adaptation of GS.
Figure 2: In the absence of a pre-existing input mesh, we parameterize Gaussians to create a structure referred to as Triangle Soup, which facilitates easy modification of the object. The Triangle Soup can be edited directly or modified using parametrization derived from estimated meshes, which do not need to be perfectly fitted. As a result, mesh-based deformation is applied to edit the object.
Recently, we have observed the emergence of several promising methods for rendering unseen views of 3D objects and scenes using neural networks. For instance, Neural Radiance Fields (NeRFs) [15] have rapidly grown in popularity within the computer vision and graphics communities [9] as they enable the creation of high-quality renders. Despite this interest and growing body of related research, the long training and inference time remains an unsolved challenge for NeRFs.
In contrast, latterly introduced Gaussian Splatting (GS) [11] offers swift training and real-time rendering capabilities. What is unique to this method is that it represents 3D objects using Gaussian distributions (i.e. Gaussians). Hence, it does not rely on any neural network. Consequently, Gaussians can be employed in a manner akin to manipulating 3D point clouds or meshes, allowing for actions like resizing or repositioning in 3D space. Nonetheless, practical challenges may arise when altering Gaussian positions, particularly in accurately tracking changes in the shape of Gaussian components, such as the ellipses. Moreover, scaling Gaussian components proves challenging when the object undergoes resizing, which is not an issue for classical meshes, as their triangle faces can be readily updated when vertex positions are adjusted.
Render Reconstruction           Modifications
@@
The above constraints may be resolved by constructing Gaussian directly on the mesh, as shown by SuGaR [10]. Unlike SuGaR, we do not have two expensive stages dedicated to producing high-quality mesh. Instead, our Gaussian Mesh Splatting (GaMeS) is designed to accommodate two distinct input scenarios. First, mesh-initialized input, when a mesh is provided as input, it may either be well-aligned (our baseline model) or poorly fitted, resembling the general form of the target object. For mismatched meshes, our one-stage training approach gradually refines the mesh alignment with the object during one-stage training. Second, mesh-free input when no mesh is available, GaMeS directly parametrizes flat Gaussians by constructing a representation inspired by triangular meshes. This results in a set of unconnected triangles called Triangle Soup [^1] Triangle Soup can be called a pseudo-mesh because it imitates the mesh of a 3D object but lacks the explicit connectivity information that defines a mesh., which serves as a mesh-like structure. In such a strategy, our Triangle Soup is not dedicated to describing the object’s surface, but we can treat Triangle Soup analogously to a conventional mesh, allowing for flexible edits and object modifications (see Fig. 1). Additionally, the GaMeS can be further parameterized by estimating an approximate mesh (see Fig. 2), providing a versatile foundation for structural adjustments and fine-grained control. The mesh-like approach can also be applied to model large scenes (see Fig. 3).
Using GaMeS, we can obtain comparable to state-of-the-art high-quality outcomes that can be attained for static scenes akin to the GS method. In summary, this work makes the following contributions:
@@
  • We introduce a Triangle Soup structure as a parametrization of flat GS for representing 3D objects, enabling efficient object editing without the need for costly meshing stages.
  • Our method relies only upon essential vector operation. Consequently, we can render scenes at a similar pace to their static counterpart.
@@ 2 Related Works
Point-based Gaussian representations have found a large number of potential application scenarios, including substituting point-cloud data [6], molecular structures modeling [4], or shape reconstruction [12]. These representations can also be utilized in applications involving shadow [17] and cloud rendering [14]. In addition, a new technique of 3D-GS was recently introduced [11]. This method couples splatting methods with point-based rendering for real-time rendering speed. The rendering quality of GS is comparable to that of Mip-NeRF [1], one of the best multilayer perceptron-based renderers.
@@
The GS surpasses NeRF in terms of both training and inference speed, distinguishing itself by not depending on a neural network for operation. Instead, GS stores essential information within its 3D Gaussian components, a feature that renders it well-suited for dynamic scene modeling [24]. Additionally, integrating GS with a dedicated 3D computer graphics engine is a straightforward process [11]. However, conditioning GS is a challenging task due to the large number of Gaussian components typically involved, which can reach hundreds of thousands.
GaussianAvatars [18] utilizes the local coordinate system to generate Gaussians corresponding to the mesh’s faces. This approach is designed explicitly for avatars and assumes the availability of a realistic (external) model for mesh fitting. However, training both the mesh and GS simultaneously is not feasible. While these solutions offer certain advantages, they do not directly combine Gaussian components with the mesh. As a result, automatic adaptation of the Gaussian parameters to changing mesh is impossible in these approaches.
The more advanced method is proposed by SuGaR [10], which introduces a regularization term in the GS cost function to encourage alignment between the Gaussians and the scene surface. SuGaR achieves this by using signed distance functions (SDF) and minimizing the difference between it and computed Gaussian values. Furthermore, the authors propose a method for extracting mesh from the GS model and introduce an optional refinement strategy that binds Gaussians to the mesh surface and jointly optimizes these Gaussians and the mesh through GS rendering.
In contrast, our work shows we can produce an editable GS model without expensive pre-processing stages like those used in SuGaR.
@@ 3 GaMeS : Mesh-Based Gaussian Splatting
Here, we provide the details of the GaMeS model, commencing with the fundamental aspects of vanilla GS. Subsequently, we elucidate how we parameterize Gaussian distributions on the mesh faces. Then, we investigate extracting a mesh face from a single Gaussian component used to re-parametrize Gaussian components to make them editable. Finally, we introduce our novel GaMeS approach.
Figure 5: The left image presents the Gaussian component constructed on the face. The right image presents how our model uses previously estimated Gaussian to construct $k$ (in Fig. $k=3$) components on the face.
Gaussian Splatting (GS)
The GS technique captures a 3D scene through an ensemble of 3D Gaussians, each defined by its position (mean), covariance matrix, opacity, and color. Additionally, spherical harmonics (SH) are employed to depict the color attributes. [8, 16]. The GS effectiveness is primarily attributed to the rendering process, which utilizes projections of Gaussian components. GS relies on a dense set of 3D Gaussian components with color and opacity:
@@
In practice, GS uses factorization $\Sigma=RSS^{T}R^{T}$. We can train GS with a flat Gaussian distribution by forcing $s_{1}=0$, where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$.
The parameters of Gaussian distributions undergo direct training through gradient optimization. To enhance adaptability to complex scenes, GS employs additional training strategies. Notably, significant Gaussian components are subdivided. If the update parameters are substantial, Gaussian components are duplicated, which also can be removed due to their low transparency. In addition, the GS training procedure is implemented in the CUDA kernel, which supports fast training and real-time rendering.
@@
In GaMeS (baseline), we placed all Gaussian components on the mesh surface. Let us examine a single triangle face with vertices $V=\{{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}\}\subset\mathbb{R}^{3}.$ We aim to parameterize the Gaussian components using vertices from the face $V$. We express the mean vector as a convex combination of vertices $V$, thus determining the Gaussian splats positions:
@@
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are trainable parameters such that $\alpha_{1}+\alpha_{2}+\alpha_{3}=1$. Through this parametrization, we consistently maintain the Gaussians positioned in the middle of face $V$.
@@
The covariance matrix can be defined as an empirical covariance calculated from three points. However, such a solution is not trivial when combined with the optimization proposed in GS. Instead, the covariance is parameterized by factorization: $\Sigma=RSS^{T}R^{T},$ where $R$ is the rotation matrix and $S$ the scaling parameters. Here, we define rotation and scaling matrices to stay in the original framework. Let start from orthonormal vectors: ${\bf r}_{1},{\bf r}_{2},{\bf r}_{3}\in\mathbb{R}^{3}$, arranged into rotation matrix $R_{V}=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}].$ Where the first vector is defined by the normal vector:
@@
Original Gaussian
Create face
Parameterized Gaussian
Figure 6: Each Gaussian corresponds to one face from the Triangle Soup. The model parametrizes the Gaussian depending on the vertices, allowing it to be edited.
Refer to caption Refer to caption
Figure 7: GaMeS parametrizes all Gaussian components by mesh vertices to lie directly on the mesh. Therefore, when we modify the mesh positions, we automatically adapt the components’ means and covariances. Consequently, we can see differences in the rendered image quality for the transforming when (a) only the Gaussian position is editable, and (b) when the Gaussian rotation and scaling are editable.
The last vector is obtained through orthonormalizing the vector with respect to the existing two vectors (a single step in the Gram–Schmidt process [3]):
@@
In consequence, we obtain a rotation matrix $R_{V}=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, which aligns with the triangular face. As scaling parameter $S$ we use $S_{V}=\mathrm{diag}(s_{1},s_{2},s_{3}),$ where $s_{1}=\varepsilon,$ $s_{2}=\|\mathrm{mean}({\bf v}_{1},{\bf v}_{2},{\bf v}_{3})-{\bf v}_{1}\|,$ and $s_{3}=\langle{\bf v}_{2},{\bf r}_{3}\rangle.$ The first scaling parameters correspond with the normal vector. Since, in our case, the Gaussians are positioned on the faces, aligning with the surface $s_{1}$ should be equal to zero. To avoid numerical problems, we fix this value as a small number (constant).
Consequently, the covariance of Gaussian distribution positioned to face is given by: $\Sigma_{V}=R_{V}S_{V}S_{V}^{T}R_{V}^{T},$ and correspond with the shape of a triangle $V$. Fig.  5 shows the process of determining Gaussians. Here the $\alpha_{j}^{i}$ refer to the $i$-th Gaussian $\mathcal{G}_{V}^{i}$ and the $j$-th vertex, respectively.
@@
This distribution method enables us to render Gaussian splats solely dependent on the position of the mesh triangles. Consequently, we can readily perform geometric transformations when dealing with Gaussians within a single (triangle) face. Moreover, triangle transformation will apply Gaussian transformation as well.
The baseline demonstrates that objects can effectively be modeled using flat Gaussians and subsequently modified through a mesh-based representation.
A key limitation of our Baseline method is its dependence on an initial 3D mesh of the scene to parametrize flat Gaussians. To address this, we propose an approach that begins with a coarse or approximate mesh, such as the FLAME model, which we demonstrate in the experiments section.
However, our model ultimately bypasses the need for any mesh by directly parameterizing Gaussians through a set of unconnected triangles—-a Triangle Soup. This Mesh–Free approach allows for greater flexibility and reduces reliance on mesh accuracy, as each triangle independently defines the Gaussian parameters, achieving direct control over the representation.
Distribution using Triangle Soup
Without any pre-existing mesh, we can automatically generate Triangle Soup structure, which can subsequently be used for object modification. We start by training GS with flat Gaussian splats, setting $s_{1}=\epsilon$. Next, we estimate a triangle face for each Gaussian component and parameterize these components using the vertices. Notably, our Triangle Soup is designed specifically to modify the GS model rather than to approximate object surfaces (see Fig. 2). After re-parametrization, we obtain the final GaMeS.
Let’s assume Gaussian is defined by mean ${\bf m}$, rotation matrix $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$ and scaling $S=\mathrm{diag}(\varepsilon,s_{2},s_{3}).$ We define three vertex of a triangle: $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}],$ where ${\bf v}_{1}={\bf m}$, ${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$, see Fig. 6.
Now we freeze the vertices of the triangle $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$ and re-parameterize the Gaussian component by defining $\hat{\bf m}$, $\hat{R}=[\hat{\bf r}_{1},\hat{\bf r}_{2},\hat{\bf r}_{3}]$ and $\hat{S}=\mathrm{diag}(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}).$ First, we put $\hat{\bf m}={\bf v}_{1}$. The first vertex of $\hat{R}$ is given by the normal vector:
@@
where $\times$ is the cross product. The second one is defined by $\hat{\bf r}_{2}=\frac{({\bf v}_{2}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\|}.$ The third one is obtained as a single step in the Gram–Schmidt process [3]:
@@
Scaling parameters can also be easily calculated as: $s_{1}=\varepsilon$, $\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|$ and $\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle$. Consequently, the covariance of Gaussian distribution positioned to a triangle is given by $\hat{\Sigma}_{V}=\hat{R}_{V}\hat{S}_{V}\hat{S}_{V}^{T}\hat{R}_{V}^{T},$ and correspond with the shape of a triangle $V$. For one $V$, we define the corresponding Gaussian component:
@@
Initial Position
GaMeS
Only Gaussians means on mesh
Figure 8: Compared to a method solely based on Gaussian averaging dependent on position, we observe that (following rotation and stretching the surface) Gaussians should dynamically adjust their scale and rotation to seamlessly conform to the change as in the case of GaMeS model.
We obtain the Gaussian component parameterized by a triangle.
GaMeS: Gaussian Mesh Splatting
The essential step in our model involves positioning Gaussians on the 3D mesh-based structure of the scene or object in question. Here, we have to consider two crucial experimental scenarios, namely, with (i.e., mesh-initialized input) and without using mesh (i.e., mesh-free input) during the training.
First, we assume that a mesh is being provided and exclusively train the Gaussian components’ centers, colors, and opacity. Here, numerous methods facilitate the generation of an initial mesh for objects, but our primary focus is not on this task. Instead, we emphasize addressing distinct issues that extend beyond the scope of initializing object meshes like EG3D [5], NGLOD [22] or NeRFMeshing [19]. Additionally, we also show the initialization of the mesh using Faces Learned with an Articulated Model and Expressions (FLAME) [13] can be used.
Second, we consider the Mesh-Free Input scenario. In the latter case, initialization of the mesh is required and achieved by utilizing our strategy with extracting Triangle Soup from pre-trained GS.
Let us consider a mesh-based structure denoted as: $\mathcal{M}=(\mathcal{V},\mathcal{E}),$ where $\mathcal{V}\subset\mathbb{R}^{3}$ represents vertices, and $\mathcal{E}\subset\mathbb{R}^{3}$ signifies edges. The mesh’s faces, denoted by:
=$\mathcal{F}_{\mathcal{M}}=\left\{\left({\bf v}^{V_{i}}_{1},{\bf v}^{V_{i}}_{2},{\bf v}^{V_{i}}_{3}\right)\right\}_{i}^{n}$
are described by sets of vertices. As previously detailed, such mesh serves as a parametrization for shaping Gaussian components.